OpenStudy (anonymous):

what is the resultant of A and B if , A= -2i -5j and B = i - 4j. Explain.

myininaya (myininaya):

what is resultant?

OpenStudy (anonymous):

I think we are supposed to add

OpenStudy (anonymous):

Just add the x and y components. -2i+i=-i -4j+(-5j)=-9j So: -i-9j

myininaya (myininaya):

i dont think i ever heard of the term resultant

OpenStudy (anonymous):

Resultant vector

myininaya (myininaya):

ok lol